# Conformal e-prediction

### v4

@@ 3 Conditional conformal e-predictors
Conformal prediction and conformal e/prediction satisfy the marginal property of validity that we have discussed so far. In this section we will discuss stronger properties of validity, including what we called object-conditional and label-conditional validity in [25, Sect. 4.6]. Suppose, e.g., that we have an algorithm for diagnosing Covid. In object-conditional validity, we divide the objects into separate categories (such as men and women) and require validity inside each of the categories, not just on average. In label-conditional validity, we require validity for each possible label, in this case separate validity for people who have Covid and those who do not. More generally, we can talk about observation-conditional validity, where the categories for which we require separate validity are defined in terms of both objects and labels.
The most standard approach to conditional conformal prediction is Mondrian conformal prediction (see, e.g., [25, Sections 4.6.7–4.6.8]), where the observation space $\mathbf{Z}$ is split into a family, often finite, of disjoint categories, and conformal p-values are computed for each category separately. For a relatively narrow group of conformal predictors, object-conditional validity has been established for overlapping categories by Jung et al. [10]. This has been generalized by Gibbs et al. [6] to what may be called “fuzzy categories”: a category becomes a nonnegative function on the object space $\mathbf{X}$, with categories in the sense of subsets of $\mathbf{X}$ corresponding to the indicator functions on $\mathbf{X}$. In this section we will see that the steps of moving from disjoint to overlapping and then to “fuzzy” categories become very simple and natural in conformal e-prediction.
@@
Validity results are easy to state for group-wise taxonomies. First let us state the result in the space domain generalizing Proposition 1.
@@
The proofs of Propositions 6 and 7 show that their assumptions (exchangeability or being IID) can be weakened, similarly to the case of Mondrian conformal prediction [25, Sect. 11.3.6]. On the other hand, the conclusion of Proposition 7 can be strengthened to give an iterated-logarithm result along the lines of Proposition 5.
@@ 4 Split conformal e-predictors
where $F_{z_{1},\dots,z_{n-c}}:\mathbf{X}\to\mathbf{P}(\mathbf{Y})$ is the neural network found from $z_{1},\dots,z_{n-c}$ as training set. (Training might take a long time, but applying the rule to a new object $x$ is typically quick.) And an example of an easily computable normalizing transformation is
@@ 5 Cross-conformal e-predictors
This follows from Proposition 9 and the arithmetic mean of e-variables being an e-variable (this is obvious and discussed in detail in [28, Sect. 3]). ∎
@@
To compare informally the outputs of cross-conformal predictors [25, Sect. 4.4] and cross-conformal e-predictors, we can use the rough transformation discussed in [28, Remark 2.3]: a p-value of $p$ roughly corresponds to an e-value of $1/p$. Under this transformation, the arithmetic average of e-values corresponds to the harmonic average of p-values, and the harmonic average is always less than or equal to the arithmetic average [8, Theorem 16]. This suggests that cross-conformal e-prediction produces better results than cross-conformal prediction does. In the opposite direction, the arithmetic average of p-values corresponds to the harmonic average of e-values, which again suggests that cross-conformal e-prediction produces better results.
@@
It is easy to combine the ideas of this section and Sect. 3 to design conditional cross-conformal e-predictors (or to combine Sections 4 and 3), but we stick to the simplest cases.
@@
Proposition 3 continues to hold for cross-conformal e-predictors and, therefore, it gives its time-wise property of validity in the online mode. However, intuitively, using cross-conformal e-prediction in the online mode defeats the purpose of cross-conformal prediction: once we process $x_{1},y_{1},\dots,x_{n},y_{n}$, we would like to apply the rule that we have found (see item 1 on p. 1) to a large number of new objects rather than just one. There are more complicated settings of “weak teachers” (along the lines of [25, Sect. 3.3]) that combine cross-conformal e-prediction with time-wise validity in useful ways, but we will not discuss them further in this paper.
@@ 6 Predictive efficiency of conformal e-predictors
We developed suitable criteria of efficiency for conformal prediction in [25, Sect. 3.1] (whose notation we will use in this section, except that $e$ will stand for e-values). The idea is that a criterion of efficiency cannot be regarded as suitable if in the limiting case of infinite training and test sets it leads to very unnatural optimal nonconformity measures. An example of such an unsuitable criterion of efficiency for conformal prediction is the use of average confidence and credibility, as defined in the first edition of [25] (analogously to our definition in Sect. 1 above) and analysed in [25, Sections 3.1.6–3.1.7]. This section proposes natural criteria of efficiency for conformal e-prediction that lead to natural nonconformity e-measures.
@@
of the log e-values for the false test labels may serve as a measure of the predictive efficiency of $f$ on the test set; we would like it to be as large as possible. Let us call it the observed log criterion of efficiency (the modifier “observed” will be discussed in Remark 15). The expression (19) is natural insofar as averaging logarithms of e-values is ubiquitous in numerous contexts: see, e.g., the five reasons given in [7, (7) and Sect. 3.1].
Let us find the optimal nonconformity e-measure for the observed log criterion in the “idealised setting” akin to the one considered in [25, Sect. 3.1.4] in the case of conformal prediction. For that it will be convenient to modify slightly the definition of a nonconformity e-measure.
@@
As in [25, Sect. 3.1.5], we assume that the object space $\mathbf{X}$ and the label space $\mathbf{Y}$ are finite (the latter simply means that we are interested in pattern recognition). We identify a probability measure $Q$ on a finite set $A$ (such as $\mathbf{Y}$) with a function mapping $a\in A$ to $Q(\{a\})$, thus often omitting the curly braces in expressions such as $Q(\{a\})$. If $Q$ is a probability measure on $\mathbf{Z}=\mathbf{X}\times\mathbf{Y}$, let $Q_{\mathbf{X}}\in\mathbf{P}(\mathbf{X})$ be its marginal probability measure on $\mathbf{X}$, and let $Q_{x}\in\mathbf{P}(\mathbf{Y})$, $x\in\mathbf{X}$, be the conditional probability measure on $\mathbf{Y}$ given $x$:
@@
By Jensen’s inequality, the max in (25) is 0, and it is attained for $\gamma_{x}=0$. This completes the proof of Proposition 14.
@@
In the terminology of [25, Sect. 3.1], the criterion of efficiency (19) is “observed” in that the efficiency of the conformal e-predictor on a test observation $z_{i}$ is measured by an expression, namely $\sum_{y\neq y_{i}}\ln e_{i}^{y}$, that depends on the observed true label $y_{i}$. A natural alternative to (19) is the prior log criterion
@@
(We have already used this nonconformity e-measure with $P_{x}$ replaced by its estimate: see (18) above.) Notice that while in conformal prediction suitable (“conditionally proper”) observed and prior criteria of efficiency lead to the same optimal conformity measures [25, Theorem 3.1], in the case of conformal e-prediction the optimal nonconformity e-measures are different, albeit typically very close for a large label space $\mathbf{Y}$.
@@ 7 Conclusion
    As set predictors, conformal predictors possess a very simple property of validity: their probability of error is bounded by a prespecified constant. This advantage is lost in conformal e-prediction (although conformal e-predictors enjoy other properties of validity).
@@
    In the online prediction protocol, conformal predictors as set predictors make errors at different steps independently. Conformal e-predictors satisfy a weakened version of this property, as discussed in Sect. 2.
@@
  • Designing conditional conformal e-predictors is much easier and using e-values adds flexibility, especially as compared with Mondrian conformal predictors; see Sect. 3.
  • Cross-conformal e-predictors are provably valid, unlike cross-conformal predictors, as discussed in Sect. 5.
@@ Acknowledgments
Many thanks to Aaditya Ramdas for pointing out deficiencies in presentation in an earlier version and to Alex Balinsky and Ilia Nouretdinov for useful discussions. This research has been partially supported by Astra Zeneca, Stena Line, and Mitie.
@@ Appendix A Law of the iterated logarithm for e-flows
where $\mathcal{F}_{n}$, $n=1,2,\dots$, is the exchangeable filtration (introduced in Sect. 2: $\mathcal{F}_{n}$ is generated by $\lbag Z_{1},\dots,Z_{n-1}\rbag,Z_{n},Z_{n+1},\dots$), and
@@
The proof of Proposition 17 uses the standard basic scheme (see, e.g., [19, Sect. 4.4]). As a first step we fix an e-flow $(\xi_{n},\mathcal{G}_{n})$ and assume, without loss of generality, that it is exact (an easy way to see that there is no loss of generality is to apply the idea of coupling [4, Sect. 7.4]).
We need an auxiliary exponential supermartingale from Stout’s proof of the law of the iterated logarithm for martingales [20, Lemma 5.4.1].
@@
Our argument given above proves the following more standard law of the iterated algorithm (which we do not need in the paper).
@@
Let $\eta_{n}$, $n\in\mathbb{Z}$ be a bounded two-sided sequence of random variables adapted to a filtration $\mathcal{G}_{n}$, $n\in\mathbb{Z}$. Suppose $\mathbb{E}(\eta_{n}\mid\mathcal{G}_{n-1})=0$ for all $n\in\mathbb{Z}$. Set
@@ Appendix B Bounding the error probability of conformal e-predictors via Markov’s inequality
The first advantage of conformal prediction mentioned in Sect. Conformal e-prediction is that conformal predictors can be used as set predictors, in which case their property of validity can be expressed as a low probability of error. In principle, this can also be done in the case of conformal e-prediction, and has been done in [1, Sect. 2], but it leads to a predictor that is not admissible in the terminology of statistical decision theory [29, Sect. 1.3]. This is the topic of this appendix. For simplicity we will assume that the object and label spaces $\mathbf{X}$ and $\mathbf{Y}$ are finite.
@@
where $f$ is a conformal p-predictor [25, Proposition 2.3]. Conformal $\epsilon$-predictors are just conformal predictors packaged as set predictors.
The BB-predictor [1, Sect. 2] associated with an admissible conformal e-predictor $f$ at significance level $\epsilon$ is defined as
$\Gamma(z_{1},\dots,z_{n},x):=\left\{y:f(z_{1},\dots,z_{n},x,y)<1/\epsilon\right\}.$
@@
Let $\epsilon\in(0,1)$. The BB-predictor associated with an admissible conformal e-predictor $f$ at significance level $\epsilon$ is $\epsilon$-valid, but it is dominated by the conformal $\epsilon$-predictor based on $f$’s nonconformity e-measure.
@@
In typical cases a BB-predictor will be inadmissible, being strictly dominated by the conformal predictor based on the same nonconformity (e-)measure. This will formalized in the next proposition, for which we need two definitions.
@@
Let $\epsilon\in(0,1)$. The BB-predictor associated with an admissible conformal e-predictor $f$ at significance level $\epsilon$ is inadmissible if $f$ is generic and $d(f)\geq 1/\epsilon$.

### v5

@@ 4 Conditional conformal e-predictors
Conformal prediction and conformal e/prediction satisfy the marginal property of validity that we have discussed so far. In this section we will discuss stronger properties of validity, including what we called object-conditional and label-conditional validity in [26, Sect. 4.6]. Suppose, e.g., that we have an algorithm for diagnosing Covid. In object-conditional validity, we divide the objects into separate categories (such as men and women) and require validity inside each of the categories, not just on average. In label-conditional validity, we require validity for each possible label, in this case separate validity for people who have Covid and those who do not. More generally, we can talk about observation-conditional validity, where the categories for which we require separate validity are defined in terms of both objects and labels.
The most standard approach to conditional conformal prediction is Mondrian conformal prediction (see, e.g., [26, Sections 4.6.7–4.6.8]), where the observation space $\mathbf{Z}$ is split into a family, often finite, of disjoint categories, and conformal p-values are computed for each category separately. For a relatively narrow group of conformal predictors, object-conditional validity has been established for overlapping categories by Jung et al. [11]. This has been generalized by Gibbs et al. [7] to what may be called “fuzzy categories”: a category becomes a nonnegative function on the object space $\mathbf{X}$, with categories in the sense of subsets of $\mathbf{X}$ corresponding to the indicator functions on $\mathbf{X}$. In this section we will see that the steps of moving from disjoint to overlapping and then to “fuzzy” categories become very simple and natural in conformal e-prediction.
@@
The geometric picture given in the previous paragraph makes the design of $\mathcal{K}$-exact conditional predictors very easy; it is just a matter of picking a point with nonnegative coordinates in a simply described affine space. Even without the requirement of being $\mathcal{K}$-exact, it is a matter of picking a point in a simply described polytope.
Validity results are easy to state for group-wise taxonomies. (Remember that group-wise taxonomies allow overlapping categories, and so cover many more applications as compared with Mondrian conformal prediction.) First let us state the result in the space domain generalizing Proposition 2.
@@
The proofs of Propositions 7 and 8 show that their assumptions (exchangeability or being IID) can be weakened, similarly to the case of Mondrian conformal prediction [26, Sect. 11.3.6]. On the other hand, the conclusion of Proposition 8 can be strengthened to give an iterated-logarithm result along the lines of Proposition 6.
@@ 5 Split conformal e-predictors
where $F_{z_{1},\dots,z_{n-c}}:\mathbf{X}\to\mathbf{P}(\mathbf{Y})$ is the neural network found from $z_{1},\dots,z_{n-c}$ as training set. (Training might take a long time, but applying the rule to a new object $x$ is typically quick.) An example of an easily computable normalizing transformation is
@@ 6 Cross-conformal e-predictors
This follows from Proposition 10 and the arithmetic mean of e-variables being an e-variable (this is obvious and discussed in detail in [29, Sect. 3]). ∎
@@
To compare informally the outputs of cross-conformal predictors [26, Sect. 4.4] and cross-conformal e-predictors, we can use the rough transformation discussed in [29, Remark 2.3]: a p-value of $p$ roughly corresponds to an e-value of $1/p$. Under this transformation, the arithmetic average of e-values corresponds to the harmonic average of p-values, and the harmonic average is always less than or equal to the arithmetic average [9, Theorem 16]. This suggests that cross-conformal e-prediction produces better results than cross-conformal prediction does. In the opposite direction, the arithmetic average of p-values corresponds to the harmonic average of e-values, which again suggests that cross-conformal e-prediction produces better results.
@@
It is easy to combine the ideas of this section and Sect. 4 to design conditional cross-conformal e-predictors (or to combine Sections 5 and 4), but we stick to the simplest cases.
@@
Proposition 4 continues to hold for cross-conformal e-predictors and, therefore, it gives its time-wise property of validity in the online mode. However, the online mode entails a massive loss of their computational efficiency. Intuitively, using cross-conformal e-prediction in the online mode defeats the purpose of cross-conformal prediction: once we process $x_{1},y_{1},\dots,x_{n},y_{n}$, we would like to apply the rule that we have found (see item 1 on p. 1) to a large number of new objects rather than just one. There are more complicated settings of “weak teachers” (along the lines of [26, Sect. 3.3]) that combine cross-conformal e-prediction with time-wise validity in useful ways, but we will not discuss them further in this paper.
@@ 7 Predictive efficiency of conformal e-predictors
We developed suitable criteria of efficiency for conformal prediction in [26, Sect. 3.1] (whose notation we will use in this section, except that $e$ will stand for e-values). The idea is that a criterion of efficiency cannot be regarded as suitable if in the limiting case of infinite training and test sets it leads to very unnatural optimal nonconformity measures. An example of such an unsuitable criterion of efficiency for conformal prediction is the use of average confidence and credibility, as defined in the first edition of [26] (analogously to our definition in Sect. 2 above) and analysed in [26, Sections 3.1.6–3.1.7] (where extremely awkward features of this criterion are discussed). This section proposes natural criteria of efficiency for conformal e-prediction that lead to natural nonconformity e-measures.
@@
of the log e-values for the false test labels may serve as a measure of the predictive efficiency of $f$ on the test set; we would like it to be as large as possible. Let us call it the observed log criterion of efficiency (the modifier “observed” will be discussed in Remark 16). The expression (22) is natural insofar as averaging logarithms of e-values is ubiquitous in numerous contexts: see, e.g., the five reasons given in [8, (7) and Sect. 3.1].
Let us find the optimal nonconformity e-measure for the observed log criterion in the “idealised setting” akin to the one considered in [26, Sect. 3.1.4] in the case of conformal prediction. For that it will be convenient to modify slightly the definition of a nonconformity e-measure.
@@
As in [26, Sect. 3.1.5], we assume that the object space $\mathbf{X}$ and the label space $\mathbf{Y}$ are finite (the latter simply means that we are interested in pattern recognition). We identify a probability measure $Q$ on a finite set $A$ (such as $\mathbf{Y}$) with a function mapping $a\in A$ to $Q(\{a\})$, thus often omitting the curly braces in expressions such as $Q(\{a\})$. If $Q$ is a probability measure on $\mathbf{Z}=\mathbf{X}\times\mathbf{Y}$, let $Q_{\mathbf{X}}\in\mathbf{P}(\mathbf{X})$ be its marginal probability measure on $\mathbf{X}$, and let $Q_{x}\in\mathbf{P}(\mathbf{Y})$, $x\in\mathbf{X}$, be the conditional probability measure on $\mathbf{Y}$ given $x$:
@@
By Jensen’s inequality, the max in (28) is 0, and it is attained for $\gamma_{x}=0$. This completes the proof of Proposition 15. ∎
@@
In the terminology of [26, Sect. 3.1], the criterion of efficiency (22) is “observed” in that the efficiency of the conformal e-predictor on a test observation $z_{i}$ is measured by an expression, namely $\sum_{y\neq y_{i}}\ln e_{i}^{y}$, that depends on the observed true label $y_{i}$. A natural alternative to (22) is the prior log criterion
@@
(We have already used this nonconformity e-measure with $P_{x}$ replaced by its estimate: see (21) above.) Notice that while in conformal prediction suitable (“conditionally proper”) observed and prior criteria of efficiency lead to the same optimal conformity measures [26, Theorem 3.1], in the case of conformal e-prediction the optimal nonconformity e-measures are different, albeit typically very close for a large label space $\mathbf{Y}$.
@@ 8 Conclusion
    As set predictors, conformal predictors possess a property of validity that is both reasonably strong and very simple: their probability of error is bounded by a prespecified constant. This advantage is lost in conformal e-prediction (the probability of error being bounded by a prespecified constant becomes a very weak property of validity for conformal e-predictors, and stronger properties of validity, such as the one given at the end of B, are somewhat less intuitive).
@@
    In the online prediction protocol, conformal predictors as set predictors make errors at different steps independently. Conformal e-predictors satisfy a weakened version of this property, as discussed in Sect. 3.
@@
  • Designing conditional conformal e-predictors is much easier and using e-values adds flexibility, especially as compared with Mondrian conformal predictors; see Sect. 4.
  • Cross-conformal e-predictors are provably valid, unlike cross-conformal predictors, as discussed in Sect. 6.
@@ Acknowledgments
Many thanks to Aaditya Ramdas for pointing out deficiencies in presentation in an earlier version and to Alex Balinsky and Ilia Nouretdinov for useful discussions. The anonymous reviewers’ comments helped greatly in improving the presentation further. This research has been partially supported by Astra Zeneca, Stena Line, and Mitie.
@@ Appendix A Law of the iterated logarithm for e-flows
where $\mathcal{F}_{n}$, $n=1,2,\dots$, is the exchangeable filtration (introduced in Sect. 3: $\mathcal{F}_{n}$ is generated by $\lbag Z_{1},\dots,Z_{n-1}\rbag,Z_{n},Z_{n+1},\dots$), and
@@
The proof of Proposition 18 uses the standard basic scheme (see, e.g., [20, Sect. 4.4]). As a first step we fix an e-flow $(\xi_{n},\mathcal{G}_{n})$ and assume, without loss of generality, that it is exact (an easy way to see that there is no loss of generality is to apply the idea of coupling [5, Sect. 7.4]).
We need an auxiliary exponential supermartingale from Stout’s proof of the law of the iterated logarithm for martingales [21, Lemma 5.4.1].
@@
Our argument given above proves the following more standard law of the iterated algorithm (which we do not need in this paper).
@@
Let $\eta_{n}$, $n\in\mathbb{Z}$, be a bounded two-sided sequence of random variables adapted to a filtration $\mathcal{G}_{n}$, $n\in\mathbb{Z}$. Suppose $\mathbb{E}(\eta_{n}\mid\mathcal{G}_{n-1})=0$ for all $n\in\mathbb{Z}$. Set
@@ Appendix B Bounding the error probability of conformal e-predictors via Markov’s inequality
The first advantage of conformal prediction mentioned in Sect. 1 is that conformal predictors can be used as set predictors, in which case their property of validity can be expressed as a low probability of error. In principle, this can also be done in the case of conformal e-prediction, and has been done in [2, Sect. 2], but it leads to a predictor that is not admissible in the terminology of statistical decision theory [30, Sect. 1.3]. This is the topic of this appendix. For simplicity we will assume that the object and label spaces $\mathbf{X}$ and $\mathbf{Y}$ are finite.
@@
where $f$ is a conformal p-predictor [26, Proposition 2.3]. Conformal $\epsilon$-predictors are just conformal predictors packaged as set predictors.
Let $\epsilon\in(0,1)$ (informally, this is our target probability of error). The BB-predictor [2, Sect. 2] associated with an admissible conformal e-predictor $f$ at significance level $1/\epsilon$ is defined as
$\Gamma(z_{1},\dots,z_{n},x):=\left\{y:f(z_{1},\dots,z_{n},x,y)<1/\epsilon\right\}.$ (40)
@@
Let $\epsilon\in(0,1)$. The BB-predictor associated with an admissible conformal e-predictor $f$ at significance level $1/\epsilon$ is $\epsilon$-valid, but it is dominated by the conformal $\epsilon$-predictor based on $f$’s nonconformity e-measure.
@@
In typical cases a BB-predictor will be inadmissible, being strictly dominated by the conformal predictor based on the same nonconformity (e-)measure. This will be formalized in the next proposition, for which we need two definitions.
@@
Let $\epsilon\in(0,1)$. The BB-predictor associated with an admissible conformal e-predictor $f$ at significance level $1/\epsilon$ is inadmissible if $f$ is generic and $d(f)\geq 1/\epsilon$.
@@
The key reason for the inadmissibility of the BB-predictor in typical situations is the strength of the validity property of conformal e-prediction. To see this strength more clearly, let us generalize the definition (40) by allowing the significance level $\alpha$ to be any positive number:
$\Gamma^{\alpha}(z_{1},\dots,z_{n},x):=\left\{y:f(z_{1},\dots,z_{n},x,y)<\alpha\right\}.$ (42)
The identity
$\mathbb{E}(E)=\int_{0}^{\infty}\mathbb{P}(E\geq\alpha)\,\mathrm{d}\alpha$
allows us to state the validity property for $f$ in terms of $\Gamma^{\alpha}$ as follows: the probability of error for $\Gamma^{\alpha}$ should integrate to at most 1 over $\alpha$. This is much stronger than requiring the probability of error for $\Gamma^{\alpha}$ to be at most $1/\alpha$ for each $\alpha$. The new requirement is joint rather than being a separate requirement for each $\alpha$.